Large pies cost $7, and small pies cost $4. If 20 pies were sold for $122, how many small pies were sold?
Answer by jorel1380(2518)

p+q=20
7p+4q=122
4p+4q=80
3p=42
p=14
q=6
14 large pies and 6 small pies were sold..